Determine the required value of the missing probability to make the distribution a discrete probability distribution. x ​P(x) 3 0.34 4 ​? 5 0.28 6 0.19 ​P(4) = nothing  ​(Type an integer or a​ decimal.)
Answer #1

In probability distribution all probabilities are greater than or equal to zero and sum of the all probabilities is 1.

We want to find, P(X = 4)

P(X = 4) = 1 - [ P(X = 3) + P(X = 5) + P(X = 6) ]

=> P(X = 4) = 1 - [ 0.34 + 0.28 + 0.19 ]

=> P(X = 4) = 1 - 0.81

=> P(X = 4) = 0.19

=> P(4) = 0.19